OpenStudy (anonymous):

integral e^(-2x)tan(e^-2x)dx please give me some hints...

sam (.sam.):

\[\huge \int\limits e^{-2 x} \left(\tan [e^{-2 x}]\right) \, dx\] Try u-sub

sam (.sam.):

u=-2x du=-2dx \[-\frac{du}{2}=dx\] \[\large =\int\limits e^u(\tan(e^u))(-\frac{du}{2})\] \[\large =-\frac{1}{2} \int\limits \left(e^u \tan (e^u)\right) \, du\]

OpenStudy (anonymous):

very detailed, thank u so much ^^

sam (.sam.):

yw :)

zepdrix (zepdrix):

Hmm wouldn't you rather set the entire exponential as your u? I mean unless you don't mind doing a second substitution.

OpenStudy (anonymous):

can i let u=e^-2x then it will be \[\int\limits -\frac{ 1 }{ 2} \tan u du\] is it right?

zepdrix (zepdrix):

Yah that looks good! :)

OpenStudy (anonymous):

tysm ^^